Home‎ > ‎Crew 161‎ > ‎

0122-DailySummaryReport

MDRS Daily Summary Report

Sol 13 1/22/2016
 
And its a wrap!
 
Executive Officer Adeniyi Adebayo Reporting
 
Mission Status: Research projects completed. Crew ready to hand over to the next team.
 
Sol Activity Summary: Today was devoted to rounding off research projects and cleaning up the HAB. SOL started with team debrief, followed by breakfast. Afterwards, crew members went about their research projects. In the afternoon, we cleaned up the HAB and prepared reports for mission support and tips for the incoming crew.
 
Look Ahead Plan: We are all looking forward to welcoming the new crew to MDRS. 
 
Anomalies in work: No anomalies to report.
 
Weather: The Martian weather was warm, with clear skies.
 
Crew Physical Status:  All crew members are hale and healthy
 
EVA: There was no EVA today.
 
Reports to be filed:
Final Commander’s Report
Daily Summary Report
Journalist Report
Biologist Report
Engineering Report
End of SIM photos
Mission Summary
Food inventory form
Astronomy report
 
Support Requested: No support is requested at this time.